OpenStudy (anonymous):

** please tell me if my answer is correct! **Ratios 1/4 is to 1 1/4 as 2 is to b. b =?

OpenStudy (anonymous):

is the answer ten?

OpenStudy (anonymous):

Yes, you are correct because:

OpenStudy (anonymous):

1/4 = 0.25 1 1/4 = 1.25 1.25/0.25 = 5 2*5 = 10

OpenStudy (anonymous):

thank u! it is almost too simple. i think i should have to do more steps or something!

OpenStudy (anonymous):

appreciate help

OpenStudy (anonymous):

No problem, anytime! :D Thanks for the medal!

OpenStudy (anonymous):

:) ditto